can you solve the next number in this sequence:
-1, 6, 25, 62,

Answer by rothauserc(4718)

x(n) = n^3 -2
:
x(5) = 5^3 -2 = 123
:
therefore, the next number in the sequence -1, 6, 25, 62 is 123
: